OpenStudy (anonymous):

Determine the slope of the line that contains the points (3,7) and (-2,3). Please explain how to do so.

OpenStudy (tkhunny):

\(Slope = \dfrac{Change\;in\;y}{Change\;in\;x} = \dfrac{Rise}{Run} = \dfrac{\Delta y}{\Delta x} = \dfrac{y_{1}-y_{2}}{x_{1}-x_{2}}\) What's the chance one of those looks orsounds familiar?

OpenStudy (precal):

use the slope formula you could also, plot the points and do rise over run